OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

Metals for anyone who can help me

OpenStudy (anonymous):

 Drawing out the question will help a lot

OpenStudy (anonymous):

the gradient (steepness) of the line can be calculated by dividing the change in y co-ordinate by the change in x co-ordinate \[gardient = \frac{ \Delta y }{ \Delta x }\]

OpenStudy (anonymous):

so gradient = -1/7

OpenStudy (anonymous):

since \[gradient =\frac{ 1-2 }{ 5- (-2) }\]\[gradient = \frac{ -1 }{ 7 }\]

OpenStudy (anonymous):

im writing all of this down in my notebook

OpenStudy (anonymous):

so now we can use the formula \[y-y1=m(x-x1)\] (where y1 and x1 are the y and x co-ordinate of the same point) \[y-1=\frac{ -1 }{ 7 } (x-5)\] (i am using the co-ordinate (5,1))

OpenStudy (anonymous):

rearranging the equation gives us \[y-1 = \frac{ -x }{ 7 } +\frac{ 5 }{ 7 }\]

OpenStudy (anonymous):

since 1 can be written as 7/7, we can replace 1 by 7/7

OpenStudy (anonymous):

\[y-\frac{ 7 }{ 7 } = \frac{ -x }{ 7 } + \frac{ 5 }{ 7 }\]\[y=\frac{ -x }{ 7 } + \frac{ 5 }{ 7 } + \frac{ 7 }{ 7 }\]\[y=\frac{ -x }{ 7 }+\frac{ 12 }{ 7 }\]

OpenStudy (anonymous):

and that is it...

OpenStudy (anonymous):

the standard form of an equation of a straight line is y=mx+c where y and x are variable m=gradient (-1/7 in our case) c=y-intercept (12/7 in our case)

OpenStudy (anonymous):

so how would you write that in standard form

OpenStudy (anonymous):

that is the standard form... y=mx+c

OpenStudy (anonymous):

y=(-1/7)x + (12/7)

OpenStudy (anonymous):

it will not allow me to put that in my answer box

OpenStudy (anonymous):

what does your answer box look like?

OpenStudy (anonymous):

try 7y = -x + 12

OpenStudy (anonymous):

box box box y= box box box box box

OpenStudy (anonymous):

y= is given

OpenStudy (anonymous):

try 7y = -x +12

OpenStudy (anonymous):

yep thank you so much for the help

OpenStudy (anonymous):

okay... have a nice day